# Matching Non-Identical Objects

### v3

@@ 5.2 Robust Image Matching
We next quantitatively evaluate the proposed method on a standard image matching task but with common corruptions. To our knowledge, we present the first evaluation of the robustness of image matching under common corruptions. We address two cases: (i) both input images are corrupted, and (ii) only one of them is corrupted. The former shows the robustness against corruption, and the latter shows the robustness against domain shift (i.e., clean image domain v.s. corrupted image domain).
Figure 6: The results with different matchers [4, 38, 27]. SuperPoint [10] is used for the keypoint detection. The introduction of our module increases the matching of eyes and noses between the cheetah and husky and decreases mismatching in the background.
Dataset.
@@
We use image pairs from the MegaDepth-1500 [56] test set as the basis for various experiments. MegaDepth [25] dataset consists of 3D reconstructed data from one million images for two popular phototourism destinations, with camera poses and depth maps computed via COLMAP [50] and refined as ground truth. Common Corruptions [20] simulates the 15 types of corruptions that occur when images are captured. The severity level of corruption is set to 5, which is the most challenging condition.
Metric.
@@
We follow the standard evaluation of image matching based on relative pose estimation (e.g., see [27]). From the matching results, the relative pose between two images is computed with RANSAC [13] as follows. First, an essential matrix representing the positional transformation between the images is calculated, which is then decomposed into a rotation and a translation. The camera pose error is computed by the maximum angular error, and the area under the cumulative error curve (AUC) at $20\tcdegree$ is calculated.
Robustness against common corruptions.
@@
The left side of Tab. 1 shows the AUC of matching when input images are both corrupted. The results of the dense matching method (i.e., LoFTR and Efficient LoFTR) are given for reference. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5% to 10% increase) with a slight increase in the runtime. The average AUC even exceeds that of Efficient LoFTR. The matching methods benefit from the proposed method, particularly when images are corrupted by noises and blurs. We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method. These corruptions introduce edgy geometric patterns, and the matching methods find correspondence between them. See Sec. B.1 in the supplementary material for more detailed comparisons.
Robustness against environmental changes.
@@
As the objects to be matched are non-identical, they can appear in largely different environments. The right side of Tab. 1 shows the AUC when only one of the input images is corrupted and the other is clean. As in the left side of Tab. 1, the proposed method boosts the LightGlue and GlueStick for most cases and the average case. Interestingly, the results are better for almost all weather categories. For brightness, it is the only one of all categories that shows a decrease in accuracy, but the drop is slight. Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions. Heatmap weighting is considered to structurally complement the areas hidden by occlusion. The proposed method makes existing matchers robust to environmental changes in general, with a slight precision reduction in clean images. See Section B.2 in the supplementary material for more detailed comparisons.
5.3 Limitations
We here present failure cases. Figure 7 shows typical cases caused by (a) heatmaps not covering a whole part of objects and (b) great structural differences. In the former case, the importance scores take large only on particular parts of the objects. Further, the large importance scores on the irrelevant object disrupt the matching. This may be resolved by using a better visual explanation model [73, 3, 1, 31, 54]. For the latter case, objects with great structural differences cannot be matched as there is little correspondence between the objects, and eventually, regions of similar colors or textures are matched. To address these cases, one may introduce the vision and language model, such as the segment anything model [24, 44] to complement the visual information with text and part-level mask information.
Figure 7: Failure cases. (a) High heatmap scores are given to only a small region of objects to be matched. (b) The structures of objects are clearly different.
@@ 6 Conclusion
In this study, we addressed a novel task of non-identical object matching, where objects to be matched are only assumed semantically similar, not necessarily identical, as in prior studies. Because of the non-identicality, two objects may be captured in different environments and domains. Thus, this task also encompasses robust matching under image corruptions and domain shifts. Our method exploits object detection and visual explanation as high-level features, thereby strongly focusing existing feature matchers on objects. Our experiments demonstrated the effectiveness of the proposed method in non-identical object matching and relative pose estimation under various image corruptions.
Acknowledgement.
Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP22H03658 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
@@ Appendix A Additional matching results
In Sec. 5.1 of the main text, we evaluated non-identical object matching by inserting our method between SuperPoint [10] and LightGlue [27]. Here, “Ours” in the results refers to the case where only our weighting module is used, without applying the class similarity adapter. The matching results for various image pair show that our method enhances matchers and provides dense and consistent correspondences. To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$). Additionally, we present more results for complex scenes containing multiple objects in a single image.
@@ A.1 Success cases
Figure A shows the matching results for the same class case. As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one. The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching. Even in the case with class discrepancy, as shown in Fig. B, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals. Figure C shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings). Figure D shows cases where multiple objects are present in an image, demonstrating that our class similarity adapter selects semantically similar object pairs and provides condensed matching between the two objects. These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
@@ A.2 Failure cases
There are some cases of failure in non-identical object matching, as shown in Fig. E. It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller) or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object. Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
Figure A: Matching results in the same class case. Image pairs show objects of the same class in ImageNet [9] or COCO [26] datasets. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] only finds a small number of correspondence, and many of them are incorrect. (b) The proposed method significantly improves the matching and performs dense and consistent correspondences.
Figure B: Matching results in the class discrepancy case. Image pairs show objects with class discrepancies in ImageNet [9] or COCO [26] datasets. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds many inconsistent and scattered matches, and it is unclear which is the correct correspondence. (b) The proposed method finds almost all matches for the corresponding parts of the animal, which largely improves the matching to consistency.
Figure C: Matching results in the domain shift case. Image pairs for matching are ImageNet [9] or COCO [26] datasets with common corruptions [20], and illustrations or sketches of a motorcycle, a violin, and a cat are generated by the Stable Diffusion [47]. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] finds correct matching, but also finds mismatches with many non-correspondence areas. (b) The proposed method reduces these mismatches and further improves to denser matching between objects.
Figure D: Matching results in complex scenes containing multiple objects. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] causes correspondences to scatter across multiple objects. (b) The proposed class similarity adapter effectively selects semantically similar object pairs and provides condensed and consistent matching between the relevant parts of the objects.
Figure E: Failure cases of non-identical object matching. (a) A combination of SuperPoint (SP) [10] and LightGlue (LG) [27] fails in matching. (b) The proposed method improves several matching but still leaves many corresponding areas unmatched.
@@ Appendix B Robust image matching details
With the same setup in Sec. 5.2 of the main text, we evaluate the effectiveness of the proposed method for LightGlue [27] and GlueStick [38] on the robustness to image corruptions and domain shift in image matching. Here, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$. The results of relative pose estimation with common corruptions [20] on one or both of the input images in the MegaDepth-1500 [56] show that the proposed method improves the robustness of LightGlue and GlueStick. In all experiments, we used the pretrained sparse matchers LightGlue [27], and GlueStick [38], and the dense matchers LoFTR [56] and EfficientLoFTR [64], downloaded from their official repositories. [^4] LightGlue: https://github.com/cvg/LightGlue
@@
B.1 Robustness against common corruptions
@@
In this experiment, common corruptions are added to both input images to demonstrate robustness against corruption. Table A shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight decrease in clean accuracy. For the Noise category, LightGlue and GlueStick with our method even surpass dense matchers in both cases.
B.2 Robustness against environmental changes

### v4

@@ 5.3 Robust Image Matching
We next quantitatively evaluate the proposed method on a standard image matching task with common corruptions. To our knowledge, image matching methods have only been evaluated clean images in literature, except for a few specialized to low-lighting [24] and thermal imaging [83]. We address two cases: (i) both input images are corrupted, and (ii) only one of them is corrupted. The former shows robustness to corruptions, while the latter shows it to domain shifts (i.e., clean image domain vs. corrupted image domain).
@@
Dataset. We used the MegaDepth-1500 [79] test set because it contains many objects, making it suitable for a fair evaluation of the non-identical object matching task. This is not the case with other standard datasets [70, 13, 4, 82, 72]. MegaDepth [35] dataset consists of 3D reconstructed data from one million images for two popular phototourism destinations, with camera poses and depth maps computed via COLMAP [71] and refined as ground truth. Common Corruptions [27] simulates the 15 types of corruptions that occur when images are captured. The severity level of corruption is set to 5, which is the most challenging condition.
@@
Metric. We follow the standard evaluation of image matching based on relative pose estimation (e.g., see [37]). From the matching results, the relative pose between two images is computed with RANSAC [20] as follows. First, an essential matrix representing the positional transformation between the images is calculated, which is then decomposed into a rotation and a translation. The camera pose error is computed by the maximum angular error, and the area under the cumulative error curve (AUC) at $20\tcdegree$ is calculated.
@@
Robustness against common corruptions. The left side of Tab. 2 shows the AUC of matching when input images are both corrupted. The results of the dense matching methods (i.e., LoFTR and Efficient LoFTR) are given for reference. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly 5 % to 10 % increase) with a slight increase in runtime. The average AUC even exceeds that of Efficient LoFTR. The matching methods benefit from the proposed method, particularly when images are corrupted by noise and blur. We observe AUC decrease in a few types of corruptions, such as Snow and Frost, by the introduction of the proposed method. These corruptions introduce edgy geometric patterns, and the matcher finds correspondence between them. See Sec. B.2 in the supplementary material for more detailed comparisons.
@@
Robustness against environmental changes. As the objects to be matched are non-identical, they can appear in largely different environments. The right side of Tab. 2 shows the AUC when only one of the input images is corrupted and the other is clean. As in the left side of Tab. 2, the proposed method boosts the LightGlue and GlueStick for most cases and the average case. Interestingly, the results are better for almost all weather categories. Only the brightness category shows a decrease in accuracy, but the drop is slight. Weather categories such as Snow and Frost add strong geometry to the overall image, causing mismatches as occlusions. Heatmap weighting is considered to structurally complement the areas hidden by occlusion. Overall, the proposed method enhances robustness to environmental changes, with only a slight drop in precision on clean images. See Sec. B.3 in the supplementary for details.
@@ 6 Conclusion
We addressed a novel task of matching non-identical objects, which are only semantically similar, not necessarily identical. Due to the non-identicality, two objects may be captured in different environments and domains. Thus, robust matching under image corruptions and domain shifts is also encompassed. Our method exploits object detection and visual explanation as high-level features, thereby focusing matching on objects. The experiments showed the effectiveness of our method in non-identical object matching using TMC, a new annotation-free metric, and relative pose estimation under various image corruptions.
Acknowledgement
Kazuhiko Kawamoto and Hiroshi Kera were supported by JSPS KAKENHI Grant Numbers JP23K24914 and JP22K17962, respectively. They were also supported by Ricoh Company, Ltd.
@@ Appendix A Additional matching results
In Sec. 5.1 of the main text, we evaluated non-identical object matching by inserting our method between SuperPoint [16] and LightGlue [37]. Here, “SEW” in the results refers to the case where only our weighting module is used without applying the Non-visual Object Pairing. The matching results for various image pairs show that our method enhances matchers and provides dense and consistent correspondences. The fine-grained correspondence can be applied to tasks such as object warping, as illustrated in Fig. A. To demonstrate this generality for various image pairs, we show more results in three cases: (i) same class (i.e., $y_{A}=y_{B},D_{A}=D_{B}$), (ii) class discrepancy (i.e., $y_{A}\neq y_{B},D_{A}=D_{B}$), (iii) domain shift (i.e., $o_{A}\equiv o_{B}$, $D_{A}\neq D_{B}$). Additionally, we present more results for complex scenes containing multiple objects in a single image.
@@ A.1 Success cases
Figure E shows the matching results for the same class case. As in the main text, The introduction of the proposed method turns sparse and inconsistent matching into dense and consistent one. The combination of SuperPoint and LightGlue with our method is robust to object color and shape differences and performs dense and consistent matching. Even in the case with class discrepancy, as shown in Fig. F, the proposed method successfully corresponds the same parts (e.g., eyes to eyes, nose to nose) between various animals. Figure G shows the case of domain shift between images, where our method reduces mismatching caused by image corruptions and image style changes (e.g., photos and drawings). Figure H shows cases where multiple objects are present in an image, demonstrating that our Non-visual Object Pairing selects semantically similar object pairs and provides condensed matching between the two objects. These results indicate that the proposed method can perform accurate matching in many cases under the challenging condition of non-identical object matching.
@@ A.2 Failure cases
There are some cases of failure in non-identical object matching, as shown in Fig. I. It is difficult to match objects that are of the same class but have significantly different shapes (e.g., the presence or absence of an airplane propeller) or objects that are similar in color or size to other parts (e.g., eyes and nose that are small, round, and black) of the object. Furthermore, image corruptions sometimes hide the texture of the objects, making it difficult to pinpoint the areas to be matched.
Figure A: Object warping by homography transformation for fine-grained and consistent matching using our method. (b) By warping the object in the right image to the object position in the left image and overlaying the two images, (c) it can be confirmed that the object positions are superimposed, showing consistent alignment.
Table A: Ablation study on the presence or absence of Non-visual Object Pairing in our method (Weighting Module: SEW, Non-visual Object Pairing: NOP). The pose accuracy (AUC) at the maximum angular error of $20\tcdegree$ of the relative pose estimation from image pairs MegaDepth-1500 [79] under common corruptions [27]. The left side shows both images corrupted with the same type of corruption, while the right side shows only one of the input images corrupted and the other as a clean image. NOP selects similar objects between images, but the MegaDepth-1500 dataset assumes that the same object appears in both images. Therefore, there is little need to select the object to be matched, resulting in minimal accuracy differences with and without NOP.
[…]
keypoint detector : SuperPoint keypoint detector : SuperPoint
LG LG+SEW LG+SEW+NOP LG LG+SEW LG+SEW+NOP
None (Clean) 80.61 78.42 78.91 80.61 78.42 78.91
Gaussian Noise 43.09 53.27 52.21 27.54 41.01 40.00
Shot Noise 43.41 53.82 52.81 32.10 42.35 41.41
Impulse Noise 44.98 50.11 48.94 35.95 43.67 42.65
Defocus Blur 32.69 48.35 47.37 18.25 32.13 31.25
Frosted Glass Blur 33.37 47.88 47.25 34.25 44.80 44.29
Motion Blur 42.12 53.63 52.95 50.07 53.10 52.46
Zoom Blur 24.40 31.21 30.39 34.67 40.04 39.36
Snow 31.51 30.33 30.40 56.24 58.15 58.00
Frost 32.24 31.71 31.83 62.20 64.87 64.74
Fog 70.99 73.49 72.41 76.96 78.40 77.37
Brightness 75.48 75.08 75.21 77.14 76.98 77.06
Contrast 39.50 38.47 38.77 43.09 45.17 45.11
Elastic Transform 54.78 66.21 65.69 64.93 68.34 67.92
Pixelate 67.94 68.24 68.12 66.80 68.43 68.37
JPEG Compression 29.54 36.97 36.47 47.41 53.82 53.45
Average 44.40 50.58 50.05 48.51 54.08 53.56
@@ Appendix B Robust image matching details
With the same setup in Sec. 5.3 of the main text, we evaluate the effectiveness of the proposed method for LightGlue [37] and GlueStick [54] on the robustness to image corruptions and domain shift in image matching. First, we present an ablation study on the presence or absence of the Non-visual Object Pairing. Second, we consider the maximum angular error at $10\tcdegree$ and $5\tcdegree$ instead of $20\tcdegree$. The results of relative pose estimation with common corruptions [27] on one or both of the input images in the MegaDepth-1500 [79] show that the proposed method improves the robustness of LightGlue and GlueStick. In all experiments, we used the pretrained sparse matchers LightGlue [37], and GlueStick [54], and the dense matchers LoFTR [79] and EfficientLoFTR [89], downloaded from their official repositories. [^4] LightGlue: https://github.com/cvg/LightGlue
@@
B.1 Abration on Non-visual Object Pairing
We conducted an ablation study to assess the impact of our Non-visual Object Pairing (NOP). Recall that the NOP assigns object pairs based on the highest cosine similarity of text embeddings (CLIP) derived from their class labels, thereby focusing the Grad-CAM-based weighting on semantically matched objects. Table B.1 presents the robustness evaluation against common corruptions applied either to both images or only one of the input images, comPairing results with and without the NOP. We find that the accuracy does not vary significantly between these settings. This implies that, on the MegaDepth-1500 dataset, where image pairs typically depict the same physical object—our Grad-CAM-based weighting already assigns appropriate focus without needing explicit object selection by the NOP. Hence, the NOP’s influence is less pronounced on standard image matching datasets with identical objects. In contrast, its benefits become more evident for non-identical object matching tasks, where semantically similar but distinct objects require a more careful pairing mechanism.
B.2 Robustness against common corruptions
@@
In this experiment, common corruptions are added to both input images to demonstrate robustness against corruption. Table D shows the results of relative pose estimation for maximum angular error of $10\tcdegree$ and $5\tcdegree$. As the results show, the proposed method makes the LightGlue and GlueStick robust against most types of common corruptions (roughly, 5% to 10% increase) with a slight decrease in clean accuracy. For the Noise category, LightGlue and GlueStick with our method even surpass dense matchers in both cases.
B.3 Robustness against environmental changes